Grade 9Trigonometry

4 g Hydrogen reacts with 9*1023 chlorine molecules and form HCI gas. If the total pressure of the system after the reaction is 700 mm, what is the partial pressure of HCI.

aindree
8 Years agoGrade 9

1 Answer

jayshree
8 Years ago
9.022*10^23 molecule
mole =9.023*23/(6.022*10^23)
        =1.498 moles

4 gm  H2
mole= mass /molar mass
       =4/2
      =2 moles
 H2  + Cl2 ------> 2HCL
2         1.489       0      (initially)
cl2 being limiting raegents reacts completely & limit the reaction
so after rection moles of h2 remains will be 2-1.498 =0.5
moles of Hcl formed is 2*1.498 =3
total moles after reaction remains =3 + 0.5= 3.5
partial pressure of HCl=total pressure * mole fraction of HCl
                                   =700 *(3/(0.5+3)
                                   =600 mm